OpenStudy (anonymous):

how do i write this number to a fraction 34.24

OpenStudy (anonymous):

rewrite as \[34+\frac{ 24 }{ 100 }\] does that help you so far?

OpenStudy (anonymous):

okay just simplify the fraction and we get \[34+\frac{ 6 }{ 25 }\] thus the number as a fraction is \[\frac{ (34*25)+6 }{ 25 }\]

OpenStudy (anonymous):

how did u get the six?

OpenStudy (anonymous):

why did the 24 become 25?

OpenStudy (anonymous):

i just devided each term in the fraction by 4 that's all